OpenStudy (anonymous):

Find the partial fraction decomposition of F(s)=(3s^2+3s+18)/(s^3+9s)

OpenStudy (anonymous):

\[\frac{ 3s^2 + 3s + 18 }{ s(s^2 + 9) } = \frac{ A }{ s } +\frac{ Bx + C }{ s^2 + 9 }\] should I keep going or is this enough?

OpenStudy (anonymous):

X?

OpenStudy (anonymous):

sorry. Bs + C

OpenStudy (anonymous):

so used to using x :P

OpenStudy (anonymous):

You'll want to solve for A, B and C

OpenStudy (anonymous):

What is the pattern for these kind of questions?

OpenStudy (anonymous):

The coefficient numerators are one polynomial order lower than the denominator. You must reduce the denominator in a multiple of the simplest polynomials. If it were s^3 -9s we'd have s(s^2 -9) = s(s-3)(s+3) and we'd be dealing with A/s + B/(s-3) + C/(s+3)

OpenStudy (anonymous):

If we had s^4 + 9s in the denominator, we'd be dealing with: \[\frac{ A }{ s } + \frac{ Bx^2 + Cx + D }{ s^3 + 9 }\]

OpenStudy (anonymous):

Or by pattern, do you mean how to solve or A, B and C?

OpenStudy (anonymous):

Thx,I get the way!

OpenStudy (anonymous):

cool :)

OpenStudy (anonymous):

Haha Bs :P . XD .